1,

Elite Trailer Parks has an operating profit of $285,000. Interest expense for the year was $30,500; preferred dividends paid were $28,900; and common dividends paid were $36,800. The tax was $68,500. The firm has 21,600 shares of common stock outstanding. a. Calculate the earnings per share and the common dividends per share for Elite Trailer Parks. (Round your answers to 2 decimal places.) Earnings per share Common dividends per share b. What was the increase in retained eamings for the year? Increase in retained earnings

2,

Stein Books Inc. sold 1,500 finance textbooks for $225 each to High Tuition University in 20X1. These books cost $190 to produce. Stein Books spent $12,300 (selling expense) to convince the university to buy its books Depreciation expense for the year was $15,300. In addition, Stein Books borrowed $101,000 on January 1, 20X1, on which the company paid 14 percent interest. Both the interest and principal of the loan were paid on December 31, 20X1. The publishing firms tax rate is 30 percent. Prepare an income statement for Stein Books. Stein Books Inc. Income Statement For the Year Ending December 31, 20X1 0) 0
Answer #1

1 a: EPS = Netincome/ Shares outstanding

= (Operating profit – Interest- Tax-Preference dividend)/ Shares

= (285000- 30500-68500-28900)/ 21600

=$7.27

Common dividend per share = common dividend/ Shares outstanding

= 36800/21600

=$1.70

B: Increase in retained earnings

=Operating profit – Interest- Tax-Preference dividend- common dividend

=285000- 30500-68500-28900- 36800

=120300

2: Income Statement for the year ended 31 December 20X1

$

Sales (1500*225)                                        337500

Less: cost of production = 1500*190       285000

Gross profit                                                  52500

Less:

Selling expenses             12300

Depreciation                    15300                 27600

Operating Income                                      24900

Less: Interest (101000*14%)                     14140

Profits before tax                                        10760

Less: Taxes 30%                                          3228

Net Income                                                 7532